OpenStudy (anonymous):

The density of helium is 0.001185 g/mL. What volume will be occupied by a 0.267 g sample of gas?

OpenStudy (anonymous):

\[\rho=\frac{m}{v} => v=\frac{m}{\rho}=\frac{0.267}{0.001185}=225.32mL\]

OpenStudy (anonymous):

V=0.267 g /0.001185 g/mL = 224.32 mL

OpenStudy (anonymous):

I know this is going to sound super stupid.. but what does the P stand for?

OpenStudy (anonymous):

\[\rho\] or d is the usual caractere to represent density

OpenStudy (anonymous):

Oooohhh... told ya it was going to be stupid! Thanks anyways... :)